# Gradient estimates of the Finslerian Allen-Cahn equation

### v1

@@
Abstract. In this manuscript, we study bounded positive solutions to the Finslerian Allen-Cahn equation. The Allen-Cahn equation is widely applied and connected to many mathematical branches. We find the Finslerian Allen-Cahn equation is also an Euler-Lagrange equation to a Liapunov entropy functional. We prove the global gradient estimates of its positive solutions on compact Finsler metric measure spaces adopting the lower bounds of the weighted Ricci curvature. Moreover, as application of a new comparison theorem developed by the author, we also get a local gradient estimates on noncompact forward complete Finsler metric measure spaces with locally finite misalignment, combining with the bounds of some non-Riemannian curvatures and the lower bound mixed weighted Ricci curvature. At last, we give a Liouville type theorem of such solutions.
Mathematics Subject Classification: 35K55, 53C60, 58J35
Keywords: Finslerian Allen-Cahn equation, Mixed weighted Ricci curvature, Gradient estimate, metric measure space
@@ 1 Introduction
Solutions to the Allen-Cahn equation have the intricate connection to the minimal surface theory. A well known question of Yau asks that “do all 3-manifolds contain infinite immersed minimal surfaces?” [34]. Almgren and Pitts [20] have developed a far-reaching theory of existence and regularity (cf. [23]) of min-max (unstable) minimal hypersurfaces. In particular, they proved that in a closed Riemannian manifold $(M^{n},g)$, with $n\geq 3$, there exists a closed minimal hypersurface which is smooth and embedded outside a singular set of Hausdorff dimension at most $n-8$. Such a minimal hypersurface is called a minimal hypersurface with optimal regularity. Guaraco introduced a new approach for the min-max construction of minimal hypersurfaces which was further developed by Gaspar and Guaraco [11], based on the study of the limiting behaviour of solutions to the Allen-Cahn equation. They showed that the Allen-Cahn widths are greater than or equal to the Almgren-Pitts widths. Later, A. Dey proved that the Almgren-Pitts widths and the Allen-Cahn widths coincide [8].
The gradient estimate method was successfully employed by S.T. Yau to study harmonic functions on complete Riemann manifolds [33]. S. Hou proved the local gradient estimates for bounded positive solutions to (1.1) on a complete noncompact Riemann manifold [10]. Later, L. Zhang established the Hessian estimates of positive solutions to (1.1) [35]. N. Dung and N. Khanh derived the gradient estimates of positive solutions to Allen-Cahn equation on weighted Riemann manifolds [14].
As a generalization of such method, the gradient estimation on Finsler metric measure spaces was first given by S. Ohta to the heat equation on compact Finsler manifolds [15]. C. Xia studied the harmonic function on both compact and forward complete Finsler manifolds by adopting the Moser’s iteration [31]. Later, Q. Xia provided the gradient estimates of positive solutions to the heat equation on forward complete Finsler manifolds [32] with the same technique of [31]. The author proposed a generic approach to derive the global and local gradient estimates on Finsler metric measure spaces for the Finslerian Schrödinger equation [25].
@@
$\Delta^{\nabla u}u+(1-u^{2})u=0,$ (1.3)
on a Finsler metric measure manifold $(M,F,\mu)$, where $\Delta^{\nabla u}$ is given in (2.16) below. We establish the gradient estimate of bounded positive solution of the Finslerian Allen-Cahn equation on both compact and forward complete Finslerian metric measure spaces following the method in [25]. Concretely, we prove that
Theorem 1.1.
Let $(M^{n},F,\mu)$ be a compact Finsler metric measure manifold with dimensional $n\geq 2$, whose weighted Ricci curvature satisfies $Ric^{N}\geq-K$, for some positive constant $K$. Assume the reversibility of $M$ has upper bound $\rho_{0}$. Suppose $u$ is a bounded positive solution to (1.3) on $M$ with $u\leq C$ where $C$ is a positive constant. (1) If $C\leq 1$, then we have
@@
In noncompact case, we have
Theorem 1.2.
Let $(M,F,\mu)$ be a complete noncompact Finsler metric measure manifold without boundary. Denote by $B(p,2R)$ the forward geodesic ball of radius $2R$ centered at $p$ on $M$. Suppose the mixed weighted Ricci curvature satisfies that ${}^{m}Ric^{N}\geq-K(2R)$ in $B(p,2R)$ with $K(2R)\geq 0$, and the misaligment $\alpha$ satisfies $\alpha\leq A(2R)$ in $B(p,2R)$. Let $u$ be a bounded positive solution to the Finslerian Allen-Cahn equation (1.3) on $M$ with $u\leq C$, where $C$ is a positive constant. (1) If $C\leq 1$, then we have
@@
in $B(p,R)$, where $C_{0},C_{1},C_{2}$ are positive constants related to $A,K_{0},N$. (2) If $C>1$, then we have
@@
Theorem 1.3.
Let $(M,F,\mu)$ be a compact Finsler metric measure manifold or forward complete noncompact Finsler metric measure manifold without boundary, whose Ricci curvature is nonnegative. If $u$ is a solution to (1.3) with $0<u\leq 1$, then $u$ is equal to 1 identically on $M$.
This manuscript is arranged as follows. In Section 2, we briefly review some Riemannian and non-Riemannian concepts on Finsler metric measure spaces. In Section 3, we introduce the Finslerian Allen-Cahn equation with its variation functional, and some regularity discussion of its solutions. Then we give the global gradient estimates of the bounded positive solutions in Section 4 with the $CD(K,N)$ condition and the local gradient estimates in Section 5 with the mixed weighted Ricci curvature condition.
@@ 2 Finsler metrics and Finsler metric measure spaces
    $g_{ij}(x,y):=\frac{1}{2}\frac{\partial F^{2}}{\partial y^{i}\partial y^{j}}(x,y).$ (2.1)
@@
$w(g_{y}(X,Y))-g_{y}(\nabla_{Z}X,Y)-$ $g_{y}(X,\nabla_{Z}Y)=2C_{y}(\nabla_{Z}y,X,Y),$
@@
$G=y^{i}\frac{\delta}{\delta x^{i}}-2G^{i}\frac{\partial}{\partial y^{i}},$ (2.2)
@@
$Ric(v):=F^{2}(v)\sum_{i=1}^{n-1}K(y,e_{i}),$
where $e_{1},\cdots,e_{n-1},\frac{v}{F(v)}$ form an orthonormal basis of $T_{x}M$ with respect to $g_{y}$. The Landsberg curvature of $(M,F)$ is given by
@@
$U(y,W)=\sum_{i=1}^{n}(D^{W}_{e_{i}}E_{i}-\hat{D}_{e_{i}}E_{i})$ (2.7)
being a vector field on the sphere bundle, where $\hat{D}$ denotes the Levi-Civita connection of the induced Riemannian metric $\hat{g}=g_{Y}$, with $Y$ denoting the geodesic extension field of $y$ in a neighborhood of $x$. (2.7) is also a curvature, which may be considered as a kind of trace of the $T$ curvature.
@@
$1\leq\rho\leq\min\{\sqrt{\kappa},\sqrt{1/\kappa^{*}}\}.$ (2.9)
@@
$\alpha=\sup_{x\in M}\alpha_{M}(x)=\sup_{x\in M}\sup_{V,W,Y\in S_{x}M}\frac{g_{V}(Y,Y)}{g_{W}(Y,Y)}.$ (2.11)
@@
So far, we have briefly introduced some Riemannian or non-Riemannian local quantities and tensors in Finsler geometry corresponding to Riemannian geometric quantities. Next, we will introduce more tensors related to the measure $d\mu$.
@@
$\nabla^{2}r(\frac{\partial}{\partial x^{i}},\frac{\partial}{\partial x^{j}})=\nabla_{j}\nabla_{i}r=\frac{\partial^{2}r}{\partial x^{i}\partial x^{j}}-\Gamma^{k}_{ij}(x,\nabla r)\frac{\partial r}{\partial x^{k}}.$ (2.12)
In (2.12), the derivative is taken in the direction $\nabla r$ naturally. Generally, we can take the derivative in any direction. Suppose $V$ is a local vector field around $x$ on $M$.
@@
A Finsler metric measure space $(M,F,d\mu)$ is a Finsler manifold equipped with an given measure $\mu$. In local coordinates $\{x^{i}\}_{i=1}^{n}$, we can express the volume form as $d\mu=\sigma(x)dx^{1}\wedge\cdots\wedge dx^{n}$. For any $y\in T_{x}M\setminus\{0\}$, define
@@
Definition 2.3 ([26][28]).
Suppose $(M,F,d\mu)$ is a Finsler metric measure space. For any point $x\in M$, let $\gamma=\gamma(t)$ be a forward geodesic from $x$ with the initial tangent vector $\dot{\gamma}(0)=y$. The S-curvature of $(M,F,d\mu)$ is
@@
Definition 2.3 means that the S-curvature is the changing of distortion along the geodesic in direction $y$. Modeled on the definition of $T$ curvature in [27], [25] defined the difference of $\nabla\tau$ on the tangent sphere, denoted by $\mathcal{T}$,
Definition 2.4 ([25]).
@@
$\mathcal{T}(V,W):=\nabla^{V}\tau(V)-\nabla^{W}\tau(W),$ (2.13)
@@
$\mathcal{T}_{i}=\frac{\delta}{\delta x^{i}}\tau_{k}(V)-\frac{\delta}{\delta x^{i}}\tau_{k}(W).$ (2.14)
@@
$div_{d\mu}V=\sum_{i=1}^{n}(\frac{\partial V^{i}}{\partial x^{i}}+V^{i}\frac{\partial\Phi}{\partial x^{i}}).$ (2.15)
@@
$\Delta_{d\mu}f:=div_{d\mu}(\nabla f).$ (2.16)
Noticing that $\Delta_{d\mu}f=\Delta_{d\mu}^{\nabla f}f$, where $\Delta_{d\mu}^{\nabla f}f:=div_{d\mu}(\nabla^{\nabla f}f)$ is in the view of weighted Laplacian with
$\nabla^{\nabla f}f:=\begin{cases}&g^{ij}(x,\nabla f)\frac{\partial f}{\partial x^{i}}\frac{\partial}{\partial x^{j}}\quad\mbox{for }x\in M_{f};\\ &0\quad\quad\quad\quad\quad\quad\quad\,\,\,\mbox{for }x\notin M_{f}.\end{cases}$ (2.17)
@@
On the other hand, the Laplacian of a function $f$ on a Riemannian manifold is the trace of the Hessian of $f$ with respect to the Riemannian metric $g$. On a Finsler metric measure space $(M,F,d\mu)$, the weighted Hessian $\tilde{H}(f)$ of a function $f$ on $M_{f}=\{x\in M:\nabla f|_{x}\neq 0\}$ is defined in [30] by
$\tilde{H}(f)(X,Y)=\nabla^{2}f(X,Y)-\frac{S(\nabla f)}{n-1}h_{\nabla f}(X,Y),$ (2.18)
where $h_{\nabla f}$ is the angular metric form in the direction $\nabla f$, given in (2.3) It is clear that $\tilde{H}(f)$ is still a symmetric bilinear form with
$\Delta f=\mathrm{tr}_{\nabla f}\tilde{H}(f).$ (2.19)
@@
$\tilde{H}^{V}(f)(X,Y)=\nabla^{2}(f)(X,X)-\frac{S(\nabla f)}{n-1}h_{V,\nabla f}(X,X),$ (2.20)
@@
$h_{V,\nabla f}(X,Y)=g_{V}(X,Y)-\frac{1}{F_{V}^{2}(\nabla f)}g_{V}(X,\nabla f)g_{V}(Y,\nabla f),$ (2.21)
@@
Definition 2.5 ([17][28]).
Given a unit vector $V\in T_{x}M$ and an positive integral number $k$, the weighted Ricci curvature is defined by
$Ric_{k}(V):=\begin{cases}&Ric(x,V)+\dot{S}(x,V)\quad\mbox{if }S(x,V)=0\mbox{ and }k=n\mbox{ or if }k=\infty;\\ &-\infty\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\mbox{if }S(x,V)\neq 0\mbox{ and if }k=n;\\ &Ric(x,V)+\dot{S}(x,V)-\frac{S^{2}(x,V)}{k-n}\quad\quad\quad\quad\quad\quad\quad\mbox{ if }n<k<\infty,\end{cases}$ (2.22)
@@
Definition 2.6 ([25]).
Let $(M,F,d\mu)$ be a Finsler metric measure space, and $V,W\in T_{x}M$ be linearly independent vectors. The weighted flag curvature $K^{k}(V;W)$ is defined by
$K^{k}(V;W):=\begin{cases}&K(V;W)+\frac{\dot{S}(V)}{(n-1)F^{2}(V)}\,\,\,\mbox{if }S(x,V)=0\mbox{ and }k=n\mbox{ or if }k=\infty;\\ &-\infty\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\,\,\mbox{if }S(x,V)\neq 0\mbox{ and if }k=n;\\ &K(V;W)+\frac{\dot{S}(V)}{(n-1)F^{2}(V)}-\frac{S^{2}(V)}{(n-1)(k-n)F^{2}(V)}\quad\quad\quad\mbox{ if }n<k<\infty,\end{cases}$ (2.23)
@@
Moreover, the mixed weighted Ricci curvature has also been defined in [25], denoted by ${}^{m}Ric^{k}$.
Definition 2.7 ([25]).
Given two unit vectors $V,W\in T_{x}M$ and a positive integral number $k$, the mixed weighted Ricci curvature ${}^{m}Ric^{k}(V,W)=^{m}Ric^{k}_{W}(V)$ is defined by
${}^{m}Ric^{k}_{W}(V):=\begin{cases}&\mathrm{tr}_{W}R_{V}(V)+\dot{S}(x,V)\quad\mbox{if }S(x,V)=0\mbox{ and }k=n\mbox{ or if }k=\infty;\\ &-\infty\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\mbox{if }S(x,V)\neq 0\mbox{ and if }k=n;\\ &\mathrm{tr}_{W}R_{V}(V)+\dot{S}(x,V)-\frac{S^{2}(x,V)}{k-n}\quad\quad\quad\quad\quad\quad\quad\mbox{ if }n<k<\infty,\end{cases}$ (2.24)
@@
$\mathfrak{ct}_{c}(r)=\begin{cases}\sqrt{c}\cot\sqrt{c}r,\quad\,\,\,\quad c>0,\\ 1/r,\quad\quad\quad\quad\quad\,\,\,\,c=0,\\ \sqrt{-c}\cot\sqrt{-c}r,\quad c<0.\end{cases}$ (2.25)
@@
Theorem 2.9 ([30]).
Let $(M,F,d\mu)$ be a Finsler metric measure space and $r=d_{F}(p,\cdot)$ be the forward distance from $p$. Suppose that for some $N>n$, the weighted flag curvature of $M$ satisfies $K_{N}\geq\frac{N-1}{n-1}c$, where $c$ is a positive constant. Then, for any vector field $X$ on $M$, the following inequality is valid whenever $r$ is smooth. Particularly, $r<\pi/\sqrt{c}$ when $c>0$.
$\tilde{H}(r)(X,X)\leq\frac{N-1}{n-1}\mathfrak{ct}_{c}(r)(g_{\nabla r}(X,X)-g_{\nabla r}(\nabla r,X)^{2}).$ (2.26)

### v2

@@
Abstract
In this manuscript, we study bounded positive solutions to the Finslerian Allen-Cahn equation. The Allen-Cahn equation is widely applied and connected to many mathematical branches. We find the Finslerian Allen-Cahn equation is also an Euler-Lagrange equation to a Liapunov entropy functional. Adopting the $\epsilon$-trick, we prove the global gradient estimates of its positive solutions on compact Finsler metric measure spaces adopting the lower bounds of the weighted Ricci curvature. Moreover, as application of a new comparison theorem developed by the author, we also get a local gradient estimates on noncompact forward complete Finsler metric measure spaces with locally finite misalignment, combining with the bounds of some non-Riemannian curvatures and the lower bound mixed weighted Ricci curvature. At last, we give a Liouville type theorem of such solutions.
keywords:
Finslerian Allen-Cahn equation, $CD(-K,N)$ condition, mixed weighted Ricci curvature, gradient estimate, metric measure space
pacs:
MSC Classification]35K55, 53C60, 58J05
@@ 1 Introduction
Solutions to the Allen-Cahn equation have the intricate connection to the minimal surface theory. A well known question of Yau asks that “do all 3-manifolds contain infinite immersed minimal surfaces?" [35]. Almgren and Pitts [21] have developed a far-reaching theory of existence and regularity (cf. [24]) of min-max (unstable) minimal hypersurfaces. In particular, they proved that in a closed Riemannian manifold $(M^{n},g)$, with $n\geq 3$, there exists a closed minimal hypersurface which is smooth and embedded outside a singular set of Hausdorff dimension at most $n-8$. Such a minimal hypersurface is called a minimal hypersurface with optimal regularity. Guaraco introduced a new approach for the min-max construction of minimal hypersurfaces which was further developed by Gaspar and Guaraco [11], based on the study of the limiting behaviour of solutions to the Allen-Cahn equation. They showed that the Allen-Cahn widths are greater than or equal to the Almgren-Pitts widths. Later, A. Dey proved that the Almgren-Pitts widths and the Allen-Cahn widths coincide [8].
The gradient estimate method was successfully employed by S.T. Yau to study harmonic functions on complete Riemann manifolds [34]. S. Hou proved the local gradient estimates for bounded positive solutions to (1) on a complete noncompact Riemann manifold [10]. Later, L. Zhang established the Hessian estimates of positive solutions to (1) [36]. N. Dung and N. Khanh derived the gradient estimates of positive solutions to Allen-Cahn equation on weighted Riemann manifolds [14].
As a generalization of such method, the gradient estimation on Finsler metric measure spaces was first given by S. Ohta to the heat equation on compact Finsler manifolds [15]. C. Xia studied the harmonic function on both compact and forward complete Finsler manifolds by adopting the Moser’s iteration [Xia2014]. Later, Q. Xia provided the gradient estimates of positive solutions to the heat equation on forward complete Finsler manifolds [33] with the same technique of [Xia2014]. The author proposed a generic approach to derive the global and local gradient estimates on Finsler metric measure spaces for the Finslerian Schrödinger equation [26].
@@
$\Delta^{\nabla u}u+(1-u^{2})u=0,$ (3)
on a Finsler metric measure space $(M,F,\mu)$, where $\Delta^{\nabla u}$ is given in (19) below. We establish the gradient estimate of bounded positive solution of the Finslerian Allen-Cahn equation on both compact and forward complete Finslerian metric measure spaces following the method in [26]. Concretely, we prove that
Theorem 1.
Let $(M^{n},F,\mu)$ be a compact Finsler $CD(-K,N)$ space, with dimension $n\geq 2$. Assume the reversibility of $M$ has upper bound $\rho_{0}$. Suppose $u$ is a bounded positive solution to (3) on $M$ with $u\leq C$ where $C$ is a positive constant.
(1) If $C\leq 1$, then we have
@@
In noncompact case, we denote the forword distance function from a fixed point $p$ by $r(x)=d(p,x)$, and the forward geodesic ball centered at $p$ with radius $R$ by $B(p,2R0)$. Then we obtain that
Theorem 2.
Let $(M,F,\mu)$ be a complete noncompact Finsler metric measure space without boundary, whose mixed weighted Ricci curvature satisfies that ${}^{m}Ric^{N}_{\nabla r}\geq-K(2R)$ in $B(p,2R)$ with $N>n$ and $K(2R)\geq 0$. Moreover, suppose the misaligment $\alpha$ of $M$ satisfies $\alpha\leq A(2R)$ in $B(p,2R)$, and suppose the non-Riemannian curvatures $U$, $\mathcal{T}$ and $\mathrm{div}C(V)=C^{ij}_{\,\,\,k|i}(V)V^{k}\frac{\delta}{\delta x^{j}}$ satisfy the norm bounds by $F(U)+F(\mathcal{T})+F^{*}(\mathrm{div}C(V))\leq K_{0}.$ Let $u$ be a bounded positive solution to the Finslerian Allen-Cahn equation (3) on $M$ with $u\leq C$, where $C$ is a positive constant.
(1) If $C\leq 1$, then we have
@@
in $B(p,R)$, where $C_{0},C_{1},C_{2}$ are positive constants related to $A,K_{0},N$.
(2) If $C>1$, then we have
@@
Theorem 3.
Let $(M,F,\mu)$ be a forward complete Finsler metric measure space without boundary, satisfying one of the following cases.
The mixed weighted Ricci curvature of $M$ is nonnegative for some $N>n$. Moreover, $M$ admits finite misalignment and the non-Riemannian curvatures $U$, $\mathcal{T}$ and $\mathrm{div}C(V)=C^{ij}_{\,\,\,k|i}(V)V^{k}\frac{\delta}{\delta x^{j}}$ satisfy the norm bounds by $F(U)+F(\mathcal{T})+F^{*}(\mathrm{div}C(V))\leq K_{0}.$
$M$ is a compact Finsler $CD(0,N)$ space.
If $u$ is a solution to (3) with $0<u\leq 1$, then $u$ is equal to 1 identically on $M$.
This manuscript is arranged as follows. In Section 2, we briefly review some Riemannian and non-Riemannian concepts on Finsler metric measure spaces. In Section 3, we introduce the Finslerian Allen-Cahn equation with its variation functional, and some regularity discussion of its solutions. Then we give the global gradient estimates of the bounded positive solutions in Section 4 with the $CD(-K,N)$ condition and the local gradient estimates in Section 5 with the mixed weighted Ricci curvature condition.
@@ 2 Finsler metrics and Finsler metric measure spaces
    $g_{ij}(x,y):=\frac{1}{2}\frac{\partial F^{2}}{\partial y^{i}\partial y^{j}}(x,y).$ (4)
@@
$Z(g_{y}(X,Y))-g_{y}(\nabla_{Z}X,Y)-$ $g_{y}(X,\nabla_{Z}Y)=2C_{y}(\nabla_{Z}y,X,Y),$
@@
$G=y^{i}\frac{\delta}{\delta x^{i}}=y^{i}\frac{\partial}{\partial x^{i}}-2G^{i}\frac{\partial}{\partial y^{i}},$ (5)
@@
$Ric(y):=F^{2}(y)\sum_{i=1}^{n-1}K(y,e_{i}),$
where $e_{1},\cdots,e_{n-1},\frac{y}{F(y)}$ form an orthonormal basis of $T_{x}M$ with respect to $g_{y}$. The Landsberg curvature of $(M,F)$ is given by
@@
$U(y,W)=\sum_{i=1}^{n}(D^{W}_{e_{i}}E_{i}-\hat{D}_{e_{i}}E_{i})$ (10)
being a vector field on the sphere bundle, where $\hat{D}$ denotes the Levi-Civita connection of the induced Riemannian metric $\hat{g}=g_{Y}$, with $Y$ denoting the geodesic extension field of $y$ in a neighborhood of $x$. (10) is also a curvature, which may be considered as a kind of trace of the $T$ curvature.
@@
$1\leq\rho\leq\min\{\sqrt{\kappa},\sqrt{1/\kappa^{*}}\}.$ (12)
@@
$\alpha=\sup_{x\in M}\alpha_{M}(x)=\sup_{x\in M}\sup_{V,W,Y\in S_{x}M}\frac{g_{V}(Y,Y)}{g_{W}(Y,Y)}.$ (14)
@@
So far, we have briefly introduced some Riemannian or non-Riemannian local quantities and tensors in Finsler geometry corresponding to Riemannian geometric quantities. Next, we will introduce more tensors related to the measure $\mu$.
@@
$\nabla^{2}r(\frac{\partial}{\partial x^{i}},\frac{\partial}{\partial x^{j}})=\nabla_{j}\nabla_{i}r=\frac{\partial^{2}r}{\partial x^{i}\partial x^{j}}-\Gamma^{k}_{ij}(x,\nabla r)\frac{\partial r}{\partial x^{k}}.$ (15)
In (15), the derivative is taken in the direction $\nabla r$ naturally. Generally, we can take the derivative in any direction. Suppose $V$ is a local vector field around $x$ on $M$.
@@
A Finsler metric measure space $(M,F,\mu)$ is a Finsler manifold equipped with an given measure $\mu$. In local coordinates $\{x^{i}\}_{i=1}^{n}$, we can express the volume form as $d\mu=\sigma(x)dx^{1}\wedge\cdots\wedge dx^{n}$ with some positive function $\sigma(x)$. For any $y\in T_{x}M\setminus\{0\}$, define
@@
Definition 3 ([27][29]).
Suppose $(M,F,\mu)$ is a Finsler metric measure space. For any point $x\in M$, let $\gamma=\gamma(t)$ be a forward geodesic from $x$ with the initial tangent vector $\dot{\gamma}(0)=y$. The S-curvature of $(M,F,\mu)$ is
@@
Definition 3 means that the S-curvature is the changing of distortion along the geodesic in direction $y$. Modeled on the definition of $T$ curvature in [28], [26] defined the difference of $\nabla\tau$ on the tangent sphere, denoted by $\mathcal{T}$,
Definition 4 ([26]).
@@
$\mathcal{T}(V,W):=\nabla^{V}\tau(V)-\nabla^{W}\tau(W),$ (16)
@@
$\mathcal{T}_{i}=\frac{\delta}{\delta x^{i}}\tau_{k}(V)-\frac{\delta}{\delta x^{i}}\tau_{k}(W).$ (17)
@@
$div_{\mu}V=\sum_{i=1}^{n}(\frac{\partial V^{i}}{\partial x^{i}}+V^{i}\frac{\partial\Phi}{\partial x^{i}}).$ (18)
@@
$\Delta_{\mu}f:=div_{\mu}(\nabla f).$ (19)
Noticing that $\Delta_{\mu}f=\Delta_{\mu}^{\nabla f}f$, where $\Delta_{\mu}^{\nabla f}f:=div_{\mu}(\nabla^{\nabla f}f)$ is in the view of weighted Laplacian with
$\nabla^{\nabla f}f:=\begin{cases}&g^{ij}(x,\nabla f)\frac{\partial f}{\partial x^{i}}\frac{\partial}{\partial x^{j}}\quad\mbox{for }x\in M_{f};\\ &0\quad\quad\quad\quad\quad\quad\quad\,\,\,\mbox{for }x\notin M_{f}.\end{cases}$ (20)
@@
On the other hand, the Laplacian of a function $f$ on a Riemannian manifold is the trace of the Hessian of $f$ with respect to the Riemannian metric $g$. On a Finsler metric measure space $(M,F,\mu)$, the weighted Hessian $\tilde{H}(f)$ of a function $f$ on $M_{f}=\{x\in M:\nabla f|_{x}\neq 0\}$ is defined in [31] by
$\tilde{H}(f)(X,Y)=\nabla^{2}f(X,Y)-\frac{S(\nabla f)}{n-1}h_{\nabla f}(X,Y),$ (21)
where $h_{\nabla f}$ is the angular metric form in the direction $\nabla f$, given in (6) It is clear that $\tilde{H}(f)$ is still a symmetric bilinear form with
$\Delta f=\mathrm{tr}_{\nabla f}\tilde{H}(f).$ (22)
@@
$\tilde{H}^{V}(f)(X,Y)=\nabla^{2}(f)(X,X)-\frac{S(\nabla f)}{n-1}h_{V,\nabla f}(X,X),$ (23)
@@
$h_{V,\nabla f}(X,Y)=g_{V}(X,Y)-\frac{1}{F_{V}^{2}(\nabla f)}g_{V}(X,\nabla f)g_{V}(Y,\nabla f),$ (24)
@@
Definition 5 ([18][29]).
Given a unit vector $V\in T_{x}M$ and an positive number $k$, the weighted Ricci curvature is defined by
$Ric_{k}(V):=\begin{cases}&Ric(x,V)+\dot{S}(x,V)\quad\mbox{if }S(x,V)=0\mbox{ and }k=n\mbox{ or if }k=\infty;\\ &-\infty\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\mbox{if }S(x,V)\neq 0\mbox{ and if }k=n;\\ &Ric(x,V)+\dot{S}(x,V)-\frac{S^{2}(x,V)}{k-n}\quad\quad\quad\quad\quad\quad\quad\mbox{ if }n<k<\infty,\end{cases}$ (25)
@@
Definition 6 ([26]).
Let $(M,F,\mu)$ be a Finsler metric measure space, and $V,W\in T_{x}M$ be linearly independent vectors. The weighted flag curvature $K^{k}(V;W)$ is defined by
$K^{k}(V;W):=\begin{cases}&K(V;W)+\frac{\dot{S}(V)}{(n-1)F^{2}(V)}\,\,\,\mbox{if }S(x,V)=0\mbox{ and }k=n\mbox{ or if }k=\infty;\\ &-\infty\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\,\,\mbox{if }S(x,V)\neq 0\mbox{ and if }k=n;\\ &K(V;W)+\frac{\dot{S}(V)}{(n-1)F^{2}(V)}-\frac{S^{2}(V)}{(n-1)(k-n)F^{2}(V)}\par\quad\quad\quad\mbox{ if }n<k<\infty,\end{cases}$ (26)
@@
Moreover, the mixed weighted Ricci curvature with respect to vector field $W$ has also been defined in [26], denoted by ${}^{m}Ric^{k}_{W}$.
Definition 7 ([26]).
Given two unit vectors $V,W\in T_{x}M$ and a positive number $k$, the mixed weighted Ricci curvature ${}^{m}Ric^{k}(V,W)=\,^{m}Ric^{k}_{W}(V)$ is defined by
${}^{m}Ric^{k}_{W}(V):=\begin{cases}&\mathrm{tr}_{W}R_{V}(V)+\dot{S}(x,V)\quad\mbox{if }S(x,V)=0\mbox{ and }k=n\mbox{ or if }k=\infty;\\ &-\infty\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\mbox{if }S(x,V)\neq 0\mbox{ and if }k=n;\\ &\mathrm{tr}_{W}R_{V}(V)+\dot{S}(x,V)-\frac{S^{2}(x,V)}{k-n}\quad\quad\quad\quad\quad\quad\quad\mbox{ if }n<k<\infty,\end{cases}$ (27)
@@
  • ii)
    We call the mixed weighted Ricci curvature bounded from below by a constant $K$, denoted by ${}^{m}Ric^{k}_{W}\geq K$, if the inequality holds that ${}^{m}Ric^{k}_{W}(V)\geq KF^{2}(V)$.
  • iii)
    We define the mixed weighted Ricci curvature ${}^{m}Ric^{k}_{W}(V)$ to be the weighted Ricci curvature $Ric^{k}(V)$ when the reference vector $W=0$.
@@
$\mathfrak{ct}_{c}(r)=\begin{cases}\sqrt{c}\cot\sqrt{c}r,\quad\,\,\,\quad c>0,\\ 1/r,\quad\quad\quad\quad\quad\,\,\,\,c=0,\\ \sqrt{-c}\coth\sqrt{-c}r,\quad c<0.\end{cases}$ (28)
@@
Theorem 4 ([31]).
Let $(M,F,\mu)$ be a Finsler metric measure space and $r=d_{F}(p,\cdot)$ be the forward distance from $p$. Suppose that for some $N>n$, the weighted flag curvature of $M$ satisfies $K_{N}\geq\frac{N-1}{n-1}c$, where $c$ is a positive constant. Then, for any vector field $X$ on $M$, the following inequality is valid whenever $r$ is smooth. Particularly, $r<\pi/\sqrt{c}$ when $c>0$.
$\tilde{H}(r)(X,X)\leq\frac{N-1}{n-1}\mathfrak{ct}_{c}(r)(g_{\nabla r}(X,X)-g_{\nabla r}(\nabla r,X)^{2}).$ (29)